# Sparse and Faithful Explanations Without Sparse Models

### 04 Mar 2024 v2

@@ 6 Discussion and Limitations
In this work, we introduce the Sparse Explanation Value (SEV) as a new way to measure and optimize interpretability in machine learning models. SEV measures decision sparsity, which focuses on how simply predictions can be explained, instead of global sparsity. Because users care about explanations for their individual predictions (rather than everyone else’s predictions), SEV better connects models to their real-world applications. Crucially, we have found (i) that many current models naturally have low SEV, (ii) that global sparsity is not necessary for decision sparsity, and (iii) that SEV can be directly optimized without sacrificing accuracy. The use of interpretable reference values and the development of optimization algorithms (Vol-Opt and All-Opt) further highlight SEV’s potential in real-world applications.
@@
There are several limitations to our approach. One is that causal relationships may exist among features, invalidating certain transitions across the SEV hypercube. This can be addressed by adapting the definition of vertices to change features in tandem, though it requires knowledge of the causal graph. Another practical consideration is choosing an appropriate reference and SEV variant. However, the flexibility in reference and variant choice actually helps tailor analyses to specific problems. For example, individuals or businesses might consider ${}^{+}$ to understand how they can make their profiles more appealing to lenders. In contrast, examining ${}^{-}$ can help lenders or decision-makers understand the fragility of their approvals and check whether the most critical factors are being used and monitored by the model. Lastly, ${}^{\circledR}$, which excludes unchangeable characteristics, offers more actionable explanations to users. However, restricting the explanation feature set may obscure biases inherent in the model as the excluded features cannot contribute to explanations.
@@
Ultimately, SEV shifts the burden for interpretability in machine learning models from prioritizing sparse models to prioritizing sparse decisions. SEV is distinct from existing black-box explanation methods, can be computed for pre-trained models, and can directly be optimized via efficient optimization algorithms. These qualities make it a strong candidate for incorporation into real-world modeling pipelines.
@@ Appendix A Datasets Description
The Adult data is based on U.S. census data including demographics, education and job information, marital status, and capital gain/loss (Dua and Graff, 2017). The target outcome is whether an individual’s salary is above \$50,000.
@@
The Diabetes dataset is derived from 10 years (1999-2008) of clinical care at 130 US hospitals and integrated delivery networks (Dua and Graff, 2017). It comprises over 50 features representing patient and hospital outcomes. The dataset includes attributes such as race, gender, age, admission type, time in hospital, the medical specialty of the admitting physician, number of lab tests performed, number of medications, etc. We take whether the patient will revisit the hospital in 2 years as a binary indicator.
@@
The German Credit data (Dua and Graff, 2017) uses financial and demographic markers (checking account status, credit history, employment/marital status, etc.) to predict whether an individual is at risk for defaulting on a loan.
@@ C.2 Detailed SEV Distributions Plots
C.3 Local Explanations Comparsion
In this section, we show how LIME-C, SHAP-C, DiCE, and ${}^{-}$ each explain the GBDT reference classifier across different datasets. These plots are analogous to those in Figure 3(b).
@@ C.3 Local Explanations Comparsion
Figure 47: The flip counts for different local explanations in different datasets(continue)
@@
Appendix D Model Scalability and Time Comparison
D.1 SEV Optimization Time Comparison
To evaluate the time consumption of our methods, we have tried different dataset sizes. The detailed information for each dataset is shown in Appendix A. To better compare the time consumption of the different methods, we have used MLP as the baseline model and compared its running time with and without All-Opt for different datasets. All the models were run 10 times and we report the mean and standard deviation in Table 22.
Dataset reference ${}^{+}$ ${}^{-}$
COMPAS $10.5\pm 0.69$ $26.9\pm 1.06$ $21.7\pm 0.03$
Adult $61.1\pm 0.15$ $164\pm 1.56$ $136\pm 1.22$
MIMIC III $78.2\pm 1.09$ $207\pm 2.71$ $167\pm 2.42$
German Credit $1.81\pm 0.01$ $5.01\pm 0.01$ $4.08\pm 0.04$
FICO $16.9\pm 0.78$ $47.6\pm 0.34$ $37.4\pm 1.30$
Diabetes $188\pm 3.08$ $606\pm 3.65$ $517\pm 3.31$
Table 22: The time (seconds) comparison with and without All-Opt in MLP
D.2 Local Explanation Time Comparsion
Here, we compare the time required to generate a local explanation for our method and various others: SHAP, LIME, and DiCE. We use GBDT as the baseline model and evaluate the median explanation runtime on queries in the test set.
Dataset TreeSHAP KernelSHAP LIME DiCE ${}^{-}$
MIMIC $1.13\pm 0.01$ $17.30\pm 0.10$ $6.10\pm 0.05$ Crashed $0.45\pm 0.03$
German
Credit
$0.53\pm 0.00$ $9.70\pm 0.17$ $7.11\pm 0.07$ $26.70\pm 0.91$ $1.37\pm 0.05$
FICO $0.56\pm 0.02$ $8.25\pm 0.27$ $6.05\pm 0.14$ $27.82\pm 0.05$ $15.75\pm 1.01^{*}$
Table 23: The median explanation runtime (in $10^{-2}$ s) for query units predicted as positive.
The table shows that ${}^{-}$ has a reasonable running time for black-box models on most datasets. We can observe that the explanation time is better than DiCE in all the datasets since it only flips from feature values to reference. To explain the value with a ($*{}^{*}$) in Table 23, the reason why the FICO Dataset takes a long time is that its average ${}^{-}$ is about 4, which means that the search in the SEV hypercube starting from the query is too deep. However, if the time consumption for the model takes too long, it means that this model cannot be sparsely explained and we need to use All-Opt or Vol-Opt methods to optimize our model. If we do not restrict the complexity of the explanations, it could take up to 2 days to produce all the explanations, however, as humans can handle only $7\pm 2$ cognitive entities at once (Miller, 1956), we set up the max depth for searching to depth=6 which also reduced the time consumption for model explanations.
@@ D.2 Local Explanation Time Comparsion
Appendix E Counterfactual Explanation Comparison between SEV and DiCE
In this section, we provide examples of SEV and DiCE explanations on the COMPAS dataset, to show how SEV can provide more interpretable explanations. Tables 25 and 26 show counterfactual explanations for two different queries. For the first, DiCE suggests that the number of juvenile crimes should go from 1 to 5 to change the prediction from ‘will recidivate’ to ‘won’t recidivate’. This is nonsensical, as the number of crimes should be positively correlated with the propensity to recidivate. In contrast, ${}^{-}$’s suggests that the number of prior crimes should go from 2 to 1, which makes sense. Similarly, for the second query, DiCE suggests that the number of juvenile crimes should go from 3 to 7 and the number of juvenile misdemeanors from 2 to 11 in order to change the prediction from ‘will recidivate’ to ‘won’t recidivate’. Again, this explanation is nonsensical. In contrast, ${}^{-}$ provides two possible explanations: (i) that the number of prior crimes should go from 3 to 1 or (ii) that the number of juvenile misdemeanors should go from 2 to 0. Again, this is an explanation that makes sense in a real-world context. Given that DiCE uses a random initialization to find a counterfactual explanation whereas ${}^{-}$ provides the minimum number of features to move to the ${}^{-}$ explanation, we would expect ${}^{-}$ to be more stable and to generate more interpretable explanations.
Explanations Sex Age
Juv fel
counts
Juv misd
count
Juvenile
crimes
Priors
counts
Current Charge
Degree
Predicted
Label
Query 1 Male 23 0 1 1 2 Not Felony 1
DiCE 5 0
${}^{-}$ 1 0
Table 25: DiCE and ${}^{-}$ explanations for an individual in COMPAS.
Explanations Sex Age
Juv fel
counts
Juv misd
count
Juvenile
crimes
Priors
counts
Current Charge
Degree
Predicted
Label
Query 2 Male 19 1 2 3 3 Felony 1
DiCE 11 7 0
${}^{-}$ 0 0
1 0
Table 26: DiCE and ${}^{-}$ explanations for another individual in COMPAS.
Appendix F The performance of ${}^{\circledR}$
@@
Figure 49: ${}^{\circledR}$ performance in Adult Dataset
@@
Appendix G Proof of Theorem 5
@@
Appendix H The difference between explanations of (Chen et al., 2018) and sparse explanation value
The work of (Chen et al., 2018) is fundamentally different from ours in several ways.
@@
  • Third, SEV can be used for many different kinds of models, whereas their approach must be applied only to differentiable models, such as the neural networks. (It cannot be used for decision trees, for example.) Because they use a complex neural network as their model, the computation within the explanation is a black box – we do not know how the features are combined to form a prediction.
@@
Appendix I Examples for ${}^{+}$ and ${}^{-}$ explanations
In this section, we show some examples of ${}^{+}$ and ${}^{-}$ explanations. Tables 27 to 30 show a few ${}^{+}$ and ${}^{-}$ calculations for various queries. For ${}^{+}$ explanations, we show the query, the reference, and values that are aligned to the query (from the reference). For ${}^{-}$ we show the values aligned to the reference (from the query). For instance, taking the first query of Table 27, the first row represents the query, and the second row is the reference. The following two rows show different ${}^{+}$ explanations: if we start from the reference and change its 9.6 year education to the query’s 13 years, and change the occupation from administrative clerical to executive manager, the income will be over 50K. Or, if, starting from the reference, education year is changed to 13, and the marital status is changed to civilian spouse, then his income will be over 50K. The next three lines are the ${}^{-}$ explanations: starting from the query, if we change either the year of education, the hours worked per week, or the marital status from the query values to the reference values, the income will be less than 50K.
@@ 13 Married-civ-spouse Wife
Table 30: Examples ${}^{+}$${}^{-}$ Explanations for German Credit Dataset. 0DM in the Checking account status column means that checking amount has no Deutschmarks (DM) in it, and 0-200DM means that the checking amount has 0-200 Deutschmarks (DM).

### 09 Mar 2024 v3

@@ 6 Discussion and Limitations
In this work, we introduce the Sparse Explanation Value (SEV) as a new way to measure and optimize interpretability in machine learning models. SEV measures decision sparsity, which focuses on how simply predictions can be explained, instead of global sparsity. Because users care about explanations for their individual predictions (rather than everyone else’s predictions), SEV better connects models to their real-world applications. Crucially, we have found (i) that many current models naturally have low SEV, (ii) that global sparsity is not necessary for decision sparsity, and (iii) that SEV can be directly optimized without sacrificing accuracy. The use of interpretable reference values and the development of optimization algorithms (Vol-Opt and All-Opt) further highlight SEV’s potential in applications.
@@
There are several limitations to our approach. One is that causal relationships may exist among features, invalidating certain transitions across the SEV hypercube. This can be addressed by adapting the definition of vertices to change features in tandem, though it requires knowledge of the causal graph. Another practical consideration is choosing an appropriate reference and SEV variant. However, the flexibility in reference and variant choice actually helps tailor analyses to specific problems. For example, individuals or businesses might consider ${}^{+}$ to understand how they can make their profiles more appealing to lenders. In contrast, examining ${}^{-}$ can help lenders understand the fragility of their approvals and can provide useful information about important variables. Lastly, ${}^{\circledR}$, which excludes unchangeable characteristics, offers more actionable explanations to users. However, restricting explanations may obscure biases inherent in the model since the excluded features cannot contribute to explanations.
@@
Ultimately, SEV shifts the burden for interpretability in machine learning models from prioritizing sparse models to prioritizing sparse decisions. SEV is distinct from existing black-box explanation methods, can be computed for pre-trained models, and can be optimized via efficient optimization algorithms. These qualities make it a candidate for real-world modeling pipelines.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes]
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes]
    2. (b)
      The license information of the assets, if applicable. [Yes]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Yes]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
@@ Appendix A Datasets Description
The Adult data is based on U.S. census data including demographics, education and job information, marital status, and capital gain/loss (Becker and Kohavi, 1996; Dua and Graff, 2017). The target outcome is whether an individual’s salary is above \$50,000.
@@
The Diabetes dataset is derived from 10 years (1999-2008) of clinical care at 130 US hospitals and integrated delivery networks (Strack et al., 2014; Dua and Graff, 2017). It comprises over 50 features representing patient and hospital outcomes. The dataset includes attributes such as race, gender, age, admission type, time in hospital, the medical specialty of the admitting physician, number of lab tests performed, number of medications, etc. We take whether the patient will revisit the hospital in 2 years as a binary indicator.
@@
The German Credit data (Hofmann, 1994; Dua and Graff, 2017) uses financial and demographic markers (checking account status, credit history, employment/marital status, etc.) to predict whether an individual is at risk for defaulting on a loan.
@@ C.2 Detailed SEV Distributions Plots
C.3 Local Explanations Comparison
In this section, we show how LIME-C, SHAP-C, DiCE, and ${}^{-}$ each explain a fixed GBDT reference classifier across different datasets. These plots are analogous to those in Figure 3(b). These figures do not include models optimized for SEV.
@@ C.3 Local Explanations Comparison
Figure 47: The flip counts for different local explanations in different datasets (continued)
@@
Appendix D Local Explanation Time Comparsion between Kernel SHAP and ${}^{-}$
Here, we compare the time required to generate a local explanation for our method and KernelSHAP. We use GBDT as the baseline model and evaluate the median explanation runtime on queries in the test set.
@@ Appendix D Local Explanation Time Comparsion between Kernel SHAP and ${}^{-}$
Appendix E The performance of ${}^{\circledR}$
@@
Figure 49: ${}^{\circledR}$ performance on Adult Dataset
@@
Appendix F Proof of Theorem 5
@@
Appendix G The difference between explanations of Chen et al. (2018) and sparse explanation value
The work of Chen et al. (2018) is fundamentally different from ours in several ways.
@@
  • Third, SEV can be used for many different kinds of models, whereas their approach must be applied only to differentiable models, such as neural networks. (It cannot be used for decision trees, for example.) Because they use a complex neural network as their model, the computation within the explanation is a black box – we do not know how the features are combined to form a prediction.
@@
Appendix H Examples for ${}^{+}$ and ${}^{-}$ explanations
In this section, we show some examples of ${}^{+}$ and ${}^{-}$ explanations. Tables 26 to 29 show a few ${}^{+}$ and ${}^{-}$ calculations for various queries. For ${}^{+}$ explanations, we show the query, the reference, and values that are aligned to the query (from the reference). For ${}^{-}$, we show the values aligned to the reference (from the query). For instance, taking the first query of Table 26, the first row represents the query, and the second row is the reference. The following two rows show different ${}^{+}$ explanations: if we start from the reference and change its 9.6 year education to the query’s 13 years, and change the occupation from administrative clerical to executive manager, the income will be predicted as over 50K. Or, if, starting from the reference, education year is changed to 13, and the marital status is changed to civilian spouse, then income will be predicted as over 50K. The next three lines are the ${}^{-}$ explanations: starting from the query, if we change either the year of education, the hours worked per week, or the marital status from the query values to the reference values, the income will be predicted as less than 50K.
@@ 13 Married-civ-spouse Wife
Table 29: Examples ${}^{+}$${}^{-}$ Explanations for German Credit Dataset. 0DM in the Checking account status column means that checking amount has no Deutsche marks (DM) in it, and 0-200DM means that the checking amount has 0-200 Deutsche marks (DM).